# CTPU-PTC-23-29

### v1

@@ Abstract
We demonstrate that the recently announced signal for a stochastic gravitational wave background (SGWB) from pulsar timing array (PTA) observations, if attributed to new physics, is compatible with primordial GW production due to axion-gauge dynamics during inflation. More specifically we find that axion-$U(1)$ models may lead to sufficient particle production to explain the signal while simultaneously source some fraction of sub-solar mass primordial black holes (PBHs) as a signature. Moreover there is a parity violation in GW sector, hence the model suggests chiral GW search as a concrete target for future. We further analyze the axion-$SU(2)$ coupling signatures and find that in the low backreaction regime, it is incapable of producing PTA evidence and violates PBH constraints.
@@
Introduction. There is a strong evidence for a stochastic gravitational wave background (SGWB) in Pulsar Timing Array (PTA) data from the NANOGRav [1, 2], EPTA [3], PPTA [4] and CPTA [5] experiments in the nHz frequency regime. While this observed signal is mainly thought to be of standard astrophysical origin sourced by supermassive black hole binary mergers [6, 7], on top of astrophysical background, there may be a possibility that the data may also imply SGWB of cosmological origin, such as cosmic strings [8, 9, 10, 11], cosmological phase transitions [12, 13, 14, 15, 16, 17, 18, 19, 20], scalar induced GW (SIGW) [21, 22, 23, 24, 25, 26].
@@
How about a possibility of primordial gravitational waves (GW) from cosmic inflation [27, 28, 29, 30]? In the standard inflationary model, the amplitude of SGWB is nearly scale-invariant and is parameterized by a ratio of tensor and scalar power spectra, called tensor-to-scalar ratio $r$. It is yet to be detected and current cosmic microwave background (CMB) experiments, Planck [31] and BICEP/Keck [32], have put an upper bound $r\lesssim 0.034$ at 95 % C.L. [33]. This CMB bound predicts the SGWB much smaller than that measured in PTA data. Explaining the data requires a blue-tilted spectrum around the nHz frequency regime [2, 34], which is not compatible with the standard inflationary scenario producing nearly scale invariant SGWB.
@@
In the simplest scenario discussed above, SGWB is produced by the vacuum fluctuation of the metric, and the tensor-to-scalar ratio is directly related to the energy scale of inflation. However, this is not necessarily true if additional matter fields could source the GW during inflation. In such scenarios, the dynamics of gauge fields coupled to axions, both Abelian ($U(1)$) [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] and non-Abelian ($SU(2)$) [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62] groups, have been well studied. Axions generically couple to gauge fields via topological Chern-Simons interactions, and this coupling violates the conformal invariance of the gauge field. Owing to this coupling, there is copious production of one circular polarization of the gauge field which in turn amplifies GW background during inflation. The resultant sourced SGWB is generically scale-dependent and its spectral shape is controlled by the time evolution of axion field. Therefore, depending on the potential, the sourced SGWB is enhanced on intermediate scales during inflation and is potentialy testable with future interferometers or PTAs [37, 44, 45, 46, 47, 52, 58, 63]. The amplified gauge field also inevitably sources scalar density modes, aside from GW, which could in turn lead to the generation of primordial black holes (PBHs) after inflation [45, 64, 65, 47, 66].
@@
In this Letter, we show the possibility that the current PTA data can be explained by the primordial GW sourced by axion-gauge dynamics during inflation. We consider two scenarios: $U(1)$ and $SU(2)$ models. The axion is a spectator field and realizes localized gauge field amplifications on intermediate scales. We discuss the possibility of generating a enhanced power spectrum of tensor modes compatible with the NANOGrav data from these models while satisfying some theoretical consistencies. We limit our analysis to the low/mild backreaction regime for which we have robust analytic expressions and remain agnostic about the dynamics in the strong backreaction regime which require approximate numerical solutions [67, 68, 69, 70, 71, 72] or full simulations on the lattice [73, 74]. Furthermore, we also discuss a possibility that the sourced power spectrum of scalar modes can lead to scalar induced GW and PBH generation.
PTA Data. Recent data from PTA experiments have found evidence for the characteristic strain GW amplitude as [1, 2]
@@
where $g_{*,0}$ is the relativistic degrees of freedom at the time of GW formation and at present, $H_{0}$ is current Hubble parameter. Then, we evaluate $\Omega_{GW,0}\sim 10^{-7}$ implying $P_{h}\sim 10^{-2}$ around frequency $yr^{-1}$, and $\Omega_{GW,0}\sim 10^{-8}$ implying $P_{h}\sim 10^{-3}$ for $f\sim(3\,yr)^{-1}$ 111 although PTAs have very-low sensitivity beyond $f>(2:3yr)^{-1}$. This is many orders of magnitude larger than current bounds on CMB scales, ie. $P_{h,CMB}<10^{-11}$. Hence the background needs to be amplified at least 9 orders of magnitude compared to vacuum fluctuations at largest scales.
@@
$\mathcal{L}_{\rm spectator}$ $=-\dfrac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi-U(\chi)-\lambda\dfrac{\chi}{4f}F_{\mu\nu}\tilde{F}^{\mu\nu}\ ,$ (4)
where $\mathcal{L}_{\rm GR}=M_{p}^{2}R/2$ and $\mathcal{L}_{\rm inflaton}=-(\partial\phi)^{2}/2-V(\phi)$ represent the Lagrangian densities of Einstein-Hilbert action and a canonical inflaton action. Regarding the inflaton potential, we let it unspecified and do not solve the background evolution of inflaton. The lagrangian density of gauge field is defined as $\mathcal{L}_{\rm gauge}=-F_{\mu\nu}F^{\mu\nu}/4$. The Hodge dual of field strength is defined as $\tilde{F}^{\mu\nu}\equiv\sqrt{-g}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$, where $\epsilon^{\mu\nu\rho\sigma}$ is an antisymmetric tensor satisfying $\epsilon^{0123}=g^{-1}$. The important thing is that the inflaton is not directly coupled to the gauge field. This assumption enables the generation of curvature perturbations sourced by the gauge field to be suppressed.
@@
$\partial_{x}^{2}A^{\pm}_{k}+\left(1\mp\dfrac{2\xi}{x}\right)A^{\pm}_{k}=0\ ,\quad\xi\equiv\dfrac{\lambda\dot{\bar{\chi}}}{2fH}\ ,$ (5)
where the dispersion relation is modified by the axion-$U(1)$ coupling controlled by a model parameter $\xi$. Initially, when the size of the mode function is deep inside the horizon ($x\gg 1$), this correction term is negligible and the gauge field obeys the standard dispersion relation. When it becomes comparable to the horizon size, however, one circular polarization mode gets an effective negative mass square for $x\lesssim 2\xi$ and a growing solution appears. The plus mode $A^{+}_{k}$ is amplified exponentially $A^{+}_{k}\propto e^{\pi\xi}$ in the time interval $(8\xi)^{-1}\lesssim x\lesssim 2\xi$. After $x\lesssim(8\xi)^{-1}$, the amplification weakens and the energy density of gauge field is diluted away by the expansion of the universe. Therefore, the gauge field production takes place at around horizon-crossing and enhances other coupled fluctuations.
This amplified gauge field enhances the coupled metric tensor modes $g_{ij}=a^{2}(\delta_{ij}+h_{ij})$ at second order level via the transverse-traceless components of energy-momentum tensor of electromagnetic field.
We define the power spectrum of GW
@@
where $P^{R(L)}_{h}$ is a dimensionless right- (left-) handed tensor power spectrum.
The shape of power spectrum is related to the time evolution of model parameter $\xi(t_{k})$, which is determined by the details of the axion potential. In this paper, we follow the previous works [43] and adopt a cosine potential
$U(\chi)=\Lambda^{4}\left[\cos\left(\dfrac{\chi}{f}\right)+1\right]\ .$ (7)
The velocity of the field gets a maximum value around $\chi(t=t_{*})=\pi f/2$. In the case of axion-$U(1)$ model, the slow-roll solutions of each inflationary model are as follows
$\xi=\dfrac{\lambda\delta}{2\cosh{[\delta H(t-t_{*})]}}\ ,\quad\delta\equiv\dfrac{\Lambda^{4}}{3H^{2}f^{2}}\ .$ (8)
@@
GW Production for U(1). In Figure 1, we plot the SGWB from axion-$U(1)$ model and compare it with that from astrophysical origin [1, 2]. The astrophysical GWB from SMBHB is expected to have $10^{-15}$ characteristic strain amplitude, and -2/3 frequency slope, which translates into $\Omega_{\rm astro}\sim 10^{-9}\left(f/yr^{-1}\right)^{2/3}$. Additionally these axion-gauge field interactions produce a unique signature, namely at small frequencies it scales as $f^{3}$ consistent with the NANOGrav slope, the peak is in log-normal shape and there is a rapid decay in the UV frequencies.
Power spectra of sourced tensor and scalar mode around their peak are parametrized as [43]
@@
where $f_{2,i}=A_{i}{\rm Exp}\left[-\ln\left(\frac{k}{x_{i}k_{*}}\right)^{2}/(2\sigma_{i}^{2})\right]$ with the amplitude $A_{i}$, the spectral width $\sigma_{i}$, and with the spectral position $x_{i}$ at which the function has the bump $k=x_{i}k_{*}$. Specifically, we have $A_{h}={\rm Exp}[-6.85+9.05\xi_{*}+0.0596\xi^{2}_{*}]$ where $\xi_{*}=\lambda\delta/2$ is the largest value of $\xi(t=t_{*})$ at the fastest motion point. We show in the shaded region the SGWB for particle production parameter in the range, $\xi_{*}=6.2$ to $\xi_{*}=6.$. We further confirm that primordial GW peak is greater than astrophysical for $\xi_{*}>5.8$. For the chosen example parameters, tensor-to-scalar ratio at the peak $r(k_{p})\sim{\rm Exp}[-0.4]\sim 1/2$. For different rolling times and particle production parameters, see [43].
A value of $P_{h}\sim 10^{-3}$ corresponds to $P_{\zeta}\sim 10^{-2.5}$ which may not violate PBH bound if fluctuations are relatively Gaussian as we will discuss in the next section. Note that although the scalar perturbations are characterized by a narrowly peaked amplitude, the GW spectrum signal is scaled as $f^{3}$ in the low frequency regime, namely in $f\ll f_{\rm peak}$. Regarding the validity of the perturbative description and effects of backreaction, we note that the values of $\xi_{*}$ explaining PTA data statistically better are close to these bounds. Beyond $\xi_{*}\gtrsim 6$ (for $\delta=0.5$), we would need a larger value of axion decay constant, closer to $M_{p}$ [75, 76]. For these parameter region, non-linear analysis becomes necessary and it is beyond the scope of our paper.
@@
Axion and SU(2) Coupling. Next, we present an inflationary model where an axion couples to an $SU(2)$ gauge field with field strength $F^{a}_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}-g\epsilon^{abc}A^{b}_{\mu}A^{c}_{\nu}$. The original model is called chromo-natural inflation, where an inflaton is directly coupled to the $SU(2)$ field [77]. However, this model was found to be inconsistent with CMB observations [78, 79, 80]222see also [81] for some challenges concerning UV completing the axion-$SU(2)$ model.. In this paper, we adopt the spectator axion-SU(2) model [54], where the Lagrangian density is given by the same form as the $U(1)$ model (3), (4).
@@
At the background level, the $SU(2)$ field can acquire an isotropic background value [82, 83] $\bar{A}^{a}_{i}(t)=a(t)Q(t)\delta^{a}_{i}$, which is diagonal between the indices of $SU(2)$ and $SO(3)$ algebra. Since this configuration respects the spatial isotropy, the background spacetime can be described by a simple FLRW metric. The axion-$SU(2)$ coupling in the presence of a non-zero isotropic gauge field vacuum expectation value, vev, induces a friction in the equation of motion for the axion. The vev is assumed to be at the bottom of its effective potential $Q(t)\simeq(-fU_{\chi}/(3g\lambda H))^{1/3}$ and the particle production is characterized by the parameter
@@
related to the $U(1)$ parameter through $\xi\simeq m_{Q}+m_{Q}^{-1}$. This solution is an attractor even if we start from the initial anisotropic parameter space as long as it leads to a stable inflationary period [84, 85].
At the level of perturbations, the fluctuation of the gauge field $\delta A^{a}_{i}$ possesses the transverse-traceless mode $\delta A^{a}_{i}\supset t^{a}_{i}$ owing to the background rotational symmetry. In the presence of the background vev, the tensor modes of the gauge field couple to GW at the linear level. In an analogous manner to the $U(1)$ case only one chirality of the tensor perturbations experiences a tachyonic instability around horizon-crossing. Hence, due to the linear coupling, only one chirality mode of GW is amplified. In this paper, we assume the amplified mode is right-handed: $t_{R}$ and $h_{R}$. The ratio of the sourced fluctuations to the vacuum fluctuations takes the form [54]
@@
For the non-Abelian case in the weak backreaction regime, it can be shown that the sourced GW can not be compatible with the NANOGrav signal. The main challenge of achieving the desired signal is due to the severe restriction imposed by the backreaction constraint. Unlike the axion-$U(1)$ case, the backreaction generally affects the EoM of the gauge field vev first (which is not present in the axion-$U(1)$ case) before there is a chance to backreact in the EoM of the axion, hence there are tighter constraints on the particle production parameter $m_{Q}$ compared to the Abelian equivalent $\xi$ [86, 75]. We extend the analysis given in detail in section 5.2 of [87]. We present our combined constraints in Figure 2 which is valid for small scales and contains the SGWB amplitude. By requiring that the sourced GW is of the order of the NANOGrav signal (e.g. $\Omega_{\rm GW}h^{2}\sim 10^{-8}$), we can produce an additional line on the same figure that transparently displays what it would take to account for the NANOGrav signal.
@@
$P_{\zeta,v,{\rm max}}\simeq\frac{g^{2}}{32\pi^{2}mQ^{4}}$ (15)
@@
Finally, we would like to point out that one can obtain an even more severe restriction in the parameter space displayed in Figure 2 by investigating whether for the NANOGrav signal at the peak $r_{\rm tot}\sim 1$ holds, which is a necessary condition for the non-overproduction of PBH. Specifically, using formulas (5.1) and (5.9) of [87] in which the sourced contributions dominate over the vacuum at the peak of the signal in the $\epsilon_{B}\gg\epsilon_{\phi}$ branch (Appendix F of the same reference), we have
@@
This ratio is always much smaller than one for $m_{Q}\gg 3.4$ which implies that the large values of $m_{Q}$ required to explain the NANOGrav signal are certain to overproduce PBH.
@@
$P_{\zeta,{\chi^{2}}}=\frac{P_{\zeta,G}^{2}}{2}\ .$ (17)
Hence, for considerable PBH abundances we have $P_{\zeta,G}\sim 2\cdot 10^{-2}$, and $P_{\zeta,\chi^{2}}\sim 10^{-3}$. For given amplitude and peak frequency, we estimate a PBH population, ranging from $10^{-10:-2}$ fraction of dark matter concentrated in masses about $10^{-4}-10^{-1}M_{\odot}$  [^4] Non-Gaussianity also affect the collapse threshold and perturbation shape, which are subdominant compared to the probability distribution function, hence neglected in this work..
@@
Enhanced scalar perturbations also source GW, scalar induced GW. In axion inflation models, the primordial GW spectrum is larger than scalar induced GW background, which is also the case in our result, for given parameters $\Omega_{{\rm SIGW,\,peak}}\sim 10^{-9}$, and leave the spectrum details to future work.
Discussion and Conclusions. The newly released PTA measurements show evidence for a SGWB. Although compatible with a background arising from supermassive black bole binary mergers, it is interesting to interpret the signal as having an early universe origin.

### v3

@@ Abstract
We demonstrate that the recently announced signal for a stochastic gravitational wave background (SGWB) from pulsar timing array (PTA) observations, if attributed to new physics, is compatible with primordial GW production due to axion-gauge dynamics during inflation. More specifically we find that axion-$U(1)$ models may lead to sufficient particle production to explain the signal while simultaneously source some fraction of sub-solar mass primordial black holes (PBHs) as a signature. Moreover there is a parity violation in GW sector, hence the model suggests chiral GW search as a concrete target for future. We further analyze the axion-$SU(2)$ coupling signatures and find that in the low/mild backreaction regime, it is incapable of producing PTA evidence and the tensor-to-scalar ratio is low at the peak, hence it overproduces scalar perturbations and PBHs.
@@
Introduction. There is a strong evidence for a stochastic gravitational wave background (SGWB) in Pulsar Timing Array (PTA) data from the NANOGRav [1, 2], EPTA [3, 4], PPTA [5] and CPTA [6] experiments in the nHz frequency regime. While this observed signal is mainly thought to be of standard astrophysical origin sourced by supermassive black hole binary mergers [7, 8, 9], on top of astrophysical background, there may be a possibility that the data also implies SGWB of cosmological origin, such as cosmic strings [10, 11, 12, 13, 14], cosmological phase transitions [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], scalar induced GW (SIGW) [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
@@
How about a possibility of primordial gravitational waves (GW) from cosmic inflation [46, 47, 48, 49]? In the standard inflationary model, the amplitude of SGWB is nearly scale-invariant and is parameterized by a ratio of tensor and scalar power spectra, called tensor-to-scalar ratio $r$. It is yet to be detected, and current cosmic microwave background (CMB) experiments, Planck [50] and BICEP/Keck [51], have put an upper bound $r\lesssim 0.034$ at 95 % C.L. [52]. This CMB bound predicts the SGWB much smaller than that measured in PTA data. Explaining the data requires a blue-tilted spectrum around the nHz frequency regime [2, 53, 54], which is not compatible with the standard inflationary scenario producing nearly scale invariant SGWB.
@@
In the simplest scenario discussed above, SGWB is produced by the vacuum fluctuation of the metric, and the tensor-to-scalar ratio is directly related to the energy scale of inflation. However, this is not necessarily true if additional matter fields could source the GW during inflation. In such scenarios, the dynamics of gauge fields coupled to axions, both Abelian ($U(1)$) [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and non-Abelian ($SU(2)$) [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] groups, have been well studied (see [84, 85] for reviews). Axions generically couple to gauge fields via topological Chern-Simons interactions, and this coupling violates the conformal invariance of the gauge field. Owing to this coupling, there is copious production of one circular polarization of the gauge field which in turn amplifies GW background during inflation. The resultant sourced SGWB is generically scale-dependent and its spectral shape is controlled by the time evolution of axion field. Therefore, depending on the potential, the sourced SGWB is enhanced on intermediate scales during inflation and is potentially testable with future interferometers or PTAs [57, 65, 66, 67, 68, 73, 79, 86]. The amplified gauge field also inevitably sources scalar density modes, aside from GW, which could in turn lead to the generation of primordial black holes (PBHs) after inflation [66, 87, 88, 68, 89] (see also PBHs via axions [90]).
@@
In this Letter, we show the possibility that the current PTA data can be explained by the primordial GW sourced by axion-gauge dynamics during inflation. We consider two scenarios: $U(1)$ and $SU(2)$ models. The axion is a spectator field and realizes localized gauge field amplifications on intermediate scales. We discuss the possibility of generating a enhanced power spectrum of tensor modes compatible with the NANOGrav data from these models while satisfying some theoretical consistencies. We limit our analysis to the low/mild backreaction regime for which we have robust analytic expressions and remain agnostic about the dynamics in the strong backreaction regime, which require approximate numerical solutions [91, 92, 93, 94, 95, 96] or full simulations on the lattice [97, 98]. We further discuss a possibility that the sourced power spectrum of scalar modes can lead to scalar induced GW and PBH generation [88, 99, 100].
PTA Data. PTA experiments found evidence for the characteristic strain amplitude as [1, 2, 4])
@@
where $g_{*}$ and $g_{0}$ are the relativistic degrees of freedom at the time of GW formation and at present, $H_{0}$ is current Hubble parameter. Then, we evaluate $\Omega_{GW,0}\sim 10^{-7}$ which implies for the tensor power spectrum that $P_{h}\sim 10^{-2}$ around frequency $yr^{-1}$, and $\Omega_{GW,0}\sim 10^{-8}$ implying $P_{h}\sim 10^{-3}$ for $f\sim(3\,yr)^{-1}$ 111 although PTAs have very-low sensitivity beyond $f>(2-3yr)^{-1}$. This is many orders of magnitude larger than current bounds on CMB scales, ie. $P_{h}(k_{CMB})<10^{-11}$, that is tensor power spectrum at CMB scales. Hence the background needs to be amplified at least 9 orders of magnitude compared to fluctuations at large scales.
@@
$\mathcal{L}_{\rm spectator}$ $=-\dfrac{1}{2}\partial_{\mu}\chi\partial^{\mu}\chi-U(\chi)-\lambda\dfrac{\chi}{4f_{a}}F_{\mu\nu}\tilde{F}^{\mu\nu}\ ,$ (4)
where $\mathcal{L}_{\rm GR}=M_{p}^{2}R/2$ and $\mathcal{L}_{\rm inflaton}=-(\partial\phi)^{2}/2-V(\phi)$ represent the Lagrangian densities of Einstein-Hilbert action and a canonical inflaton action. Regarding the inflaton potential, we let it unspecified and do not solve the background evolution of inflaton. The lagrangian density of gauge field is defined as $\mathcal{L}_{\rm gauge}=-F_{\mu\nu}F^{\mu\nu}/4$. The Hodge dual of field strength is defined as $\tilde{F}^{\mu\nu}\equiv\sqrt{-g}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/2$, where $\epsilon^{\mu\nu\rho\sigma}$ is an antisymmetric tensor satisfying $\epsilon^{0123}=g^{-1}$. The important thing is that the inflaton is not directly coupled to the gauge field  [^2] Axion being inflaton with a polynomial potential cannot generate PTA signal due to bounds on the primordial non-Gaussianity at CMB scales [101], CMB scales requires slower axion and this prevents enough enhancement at PTA scales, a way out by chaning slope around PTA scales is discussed in [66].. This assumption enables the generation of curvature perturbations sourced by the gauge field to be suppressed.
@@
$\partial_{x}^{2}A^{\pm}_{k}+\left(1\mp\dfrac{2\xi}{x}\right)A^{\pm}_{k}=0\ ,\quad\xi\equiv\dfrac{\lambda\dot{\bar{\chi}}}{2f_{a}H}\ ,$ (5)
where the dispersion relation is modified by the axion-$U(1)$ coupling controlled by a model parameter $\xi$. Initially, when the size of the mode function is deep inside the horizon ($x\gg 1$), this correction term is negligible and the gauge field obeys the standard dispersion relation. When it becomes comparable to the horizon size, however, one circular polarization mode gets an effective negative mass square for $x\lesssim 2\xi$ and a growing solution appears. The plus mode $A^{+}_{k}$ is amplified exponentially $A^{+}_{k}\propto e^{\pi\xi}$ in the time interval $(8\xi)^{-1}\lesssim x\lesssim 2\xi$. After $x\lesssim(8\xi)^{-1}$, the amplification weakens and the energy density of gauge field is diluted away by the expansion of the universe  [^3] We note that $x\lesssim(8\xi)^{-1}$ is chosen as a typical scale for the integration limit where the production of gauge fields are stronger than dilution. This scale can be chosen differently and results will be the same. Therefore, the gauge field production takes place at around horizon-crossing and enhances other coupled fluctuations.
This amplified gauge field enhances the coupled metric tensor modes $g_{ij}=a^{2}(\delta_{ij}+h_{ij})$ at second order level via the transverse-traceless components of energy-momentum tensor of electromagnetic field. We define the power spectrum of GW
@@
where $P^{R(L)}_{h}$ is a dimensionless right- (left-) handed tensor power spectrum. The shape of power spectrum is related to the time evolution of model parameter $\xi(t_{k})$, which is determined by the details of the axion potential. In this paper, we follow the previous works [64] and adopt a cosine potential
$U(\chi)=\Lambda^{4}\left[\cos\left(\dfrac{\chi}{f_{a}}\right)+1\right]\ .$ (7)
The velocity of the field gets a maximum value around $\chi(t=t_{*})=\pi f_{a}/2$, where $t_{*}$ is the moment of maximum velocity. In the case of axion-$U(1)$ model, the slow-roll solution is given as
$\xi=\dfrac{\lambda\delta}{2\cosh{[\delta H(t-t_{*})]}}\ ,\quad\delta\equiv\dfrac{\Lambda^{4}}{3H^{2}f_{a}^{2}}\ .$ (8)
@@
GW Production for U(1). In Figure 1, we plot the SGWB from axion-$U(1)$ model and compare it with that from astrophysical origin [1, 2]. The astrophysical GWB from SMBHB is expected to have $10^{-15}$ characteristic strain amplitude, and -2/3 frequency slope, which translates into $\Omega_{\rm astro}\sim 10^{-9}\left(f/yr^{-1}\right)^{2/3}$. Additionally, these axion-gauge field interactions produce a unique signature, namely at small frequencies it scales as $f^{3}$ consistent with the NANOGrav slope, the peak is in log-normal shape and there is a rapid decay in the UV frequencies.
Power spectra, ie. two-point function, of sourced tensor and scalar modes, $P_{i,s}$, around their peak are parametrized as [64]
@@
where $f_{2,i}=A_{i}{\rm Exp}\left[-\ln\left(\frac{k}{x_{i}k_{*}}\right)^{2}/(2\sigma_{i}^{2})\right]$ with the amplitude $A_{i}$, the spectral width $\sigma_{i}$, and with the spectral position $x_{i}$ at which the function has the bump $k=x_{i}k_{*}$. Specifically, we have $A_{h}={\rm Exp}[-6.85+9.05\xi_{*}+0.0596\xi^{2}_{*}]$ where $\xi_{*}=\lambda\delta/2$ is the largest value of $\xi(t=t_{*})$ at the fastest motion point. We show in the shaded region the SGWB for particle production parameter in the range, $\xi_{*}=6.2$ to $\xi_{*}=6$. We further confirm that primordial GW peak is greater than astrophysical for $\xi_{*}>5.8$. For the chosen example parameters, tensor-to-scalar ratio at the peak $r(k_{p})\sim{\rm Exp}[-0.4]\sim 1/2$. For different rolling times and particle production parameters, see [64].
A value of $P_{h}\sim 10^{-3}$  [^4] We note that due to the decay of vector modes during inflation, $N_{eff}$ bounds are only relevant for the GW background, not for the gauge field that forms $0.5-5\%$ of total energy density at its peak, which still satsifies $N_{eff}$ bounds with tiny margin. corresponds to $P_{\zeta}\sim 10^{-2.5}$ which may not violate PBH bound if fluctuations are relatively Gaussian as we will discuss in the next section. Note that although the scalar perturbations are characterized by a narrowly peaked amplitude, the GW spectrum signal is scaled as $f^{3}$ in the low frequency regime, namely in $f\ll f_{\rm peak}$.
By imposing backreaction and perturbativity bounds from the semi-analytic calculation, for $\delta=0.5$, we obtain $\xi_{*}$ = 6.11 is the maximum value that can satisfy $f_{a}<M_{p}$ by using the results in [102, 103], that is, for smaller $\xi_{*}$, the decay constant of the axion can be chosen smaller than $M_{p}$. $\xi_{*}$ larger than this value requires the decay constant of the axion to be larger than $M_{p}$, which is not viable. For this parameter region, non-linear analysis becomes necessary and it is beyond the scope of our paper.
@@
Axion and SU(2) Coupling. Next, we present an inflationary model where an axion couples to an $SU(2)$ gauge field with field strength $F^{a}_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}-g\epsilon^{abc}A^{b}_{\mu}A^{c}_{\nu}$. The original model is called chromo-natural inflation, where an inflaton is directly coupled to the $SU(2)$ field [104]. However, this model was found to be inconsistent with CMB observations [105, 106, 107]555see also [108] for some challenges concerning UV completing the axion-$SU(2)$ model.. In this paper, we adopt the spectator axion-SU(2) model [75], where the Lagrangian density is given by the same form as the $U(1)$ model (3), (4).
@@
At the background level, the $SU(2)$ field can acquire an isotropic background value [109, 110] $\bar{A}^{a}_{i}(t)=a(t)Q(t)\delta^{a}_{i}$, which is diagonal between the indices of $SU(2)$ and $SO(3)$ algebra. Since this configuration respects the spatial isotropy, the background spacetime can be described by a simple FLRW metric. The axion-$SU(2)$ coupling in the presence of a non-zero isotropic gauge field vacuum expectation value, vev, induces a friction in the equation of motion for the axion. The vev is assumed to be at the bottom of its effective potential $Q(t)\simeq(-f_{a}U_{\chi}/(3g\lambda H))^{1/3}$ and the particle production is characterized by the parameter
@@
related to the $U(1)$ parameter through $\xi\simeq m_{Q}+m_{Q}^{-1}$. This solution is an attractor even if we start from the initial anisotropic parameter space as long as it leads to a stable inflationary period [111, 112, 113].
At the level of perturbations, the fluctuation of the gauge field $\delta A^{a}_{i}$ possesses the transverse-traceless mode $\delta A^{a}_{i}\supset t^{a}_{i}$ owing to the background rotational symmetry. In the presence of the background vev, the tensor modes of the gauge field couple to GW at the linear level. In an analogous manner to the $U(1)$ case only one chirality of the tensor perturbations experiences a tachyonic instability around horizon-crossing. Hence, due to the linear coupling, only one chirality mode of GW is amplified. In this paper, we assume the amplified mode is right-handed: $t_{R}$ and $h_{R}$. The ratio of the sourced fluctuations, $P_{h,s}$, to the vacuum fluctuations, $P_{h,v}$, takes the form [75]
@@
For the non-Abelian case in the weak backreaction regime, it can be shown that the sourced GW can not be compatible with the NANOGrav signal. The main challenge of achieving the desired signal is due to the severe restriction imposed by the backreaction constraint. Unlike the axion-$U(1)$ case, the backreaction generally affects the EoM of the gauge field vev first (which is not present in the axion-$U(1)$ case) before there is a chance to backreact in the EoM of the axion, hence there are tighter constraints on the particle production parameter $m_{Q}$ compared to the Abelian equivalent $\xi$ [114, 102]. We extend the analysis given in detail in section 5.2 of [115]. We present our combined constraints in Figure 2 which is valid for small scales and contains the SGWB amplitude. We show two SGWB background values with red dashed-dotted lines, one is $\Omega_{\rm GW}h^{2}\sim 3\cdot 10^{-15}$ that is the amplitude one gets in the weak backreaction regime and the scalar vacuum perturbations are taken at the face value of the CMB scales, the other one is $\Omega_{\rm GW}h^{2}\sim 10^{-8}$, a GW signal of the order of the NANOGrav signal.
@@
$P_{\zeta,v,{\rm max}}\simeq\frac{g^{2}}{32\pi^{2}m_{Q}^{4}}$ (15)
@@
Finally, we would like to point out that one can obtain an even more severe restriction in the parameter space displayed in Figure 2 by investigating whether for the NANOGrav signal, at the peak the total amount of tensor-to-scalar ratio $r_{\rm tot}=\frac{P_{h,s}+P_{h,v}}{P_{\zeta,s}+P_{\zeta,v}}\sim\frac{P_{h,s}}{P_{\zeta,s}}\sim 1$ holds, which is a necessary condition for the non-overproduction of PBH. Specifically, using formulas (5.1) and (5.9) of [115] in which the sourced contributions dominate over the vacuum at the peak of the signal in the $\epsilon_{B}>\epsilon_{\phi}$ branch (Appendix F of the same reference), we have
@@
This ratio is always much smaller than one for $m_{Q}>3.4$ which implies that the large values of $m_{Q}$ required to explain the NANOGrav signal are certain to overproduce scalar density perturbations, and expectedly PBH.
@@
$P_{\zeta,{\chi^{2}}}\sim\frac{2}{\zeta_{c}^{2}}\;P_{\zeta,G}^{2}\ .$ (17)
$\zeta_{c}\sim\frac{9}{4}\delta_{c}$ is the critical threshold for curvature perturbations. Hence, for considerable PBH abundances we have $P_{\zeta,G}\sim 2\cdot 10^{-2}$, and $P_{\zeta,\chi^{2}}\sim 10^{-3}$. For given amplitude and peak frequency, we estimate a PBH population, ranging from $10^{-10}-10^{-2}$ fraction of dark matter concentrated in masses about $10^{-4}-10^{-1}M_{\odot}$  [^7] Non-Gaussianity also affect the collapse threshold and perturbation shape, which are subdominant compared to the probability distribution function, hence neglected in this work..
@@
Enhanced scalar perturbations also source GW, scalar induced GW. In axion inflation models, the primordial GW spectrum is larger than the scalar induced GW background, which is also the case in our result, for given parameters $\Omega_{{\rm SIGW,\,peak}}\sim 10^{-9}$ as in [100, 99, 88], and we leave the spectrum details to future work.
Discussion and Conclusions. The newly released PTA measurements show evidence of a SGWB. Although compatible with a background arising from supermassive black bole binary mergers, it is interesting to interpret the signal as having an early universe origin.